# Resolving the Ambiguity of Complete-to-Partial Point Cloud Registration for Image-Guided Liver Surgery with Patches-to-Partial Matching

### v1

@@ Abstract
In image-guided liver surgery, the initial rigid alignment between preoperative and intraoperative data, often represented as point clouds, is crucial for providing sub-surface information from preoperative CT/MRI images to the surgeon during the procedure. Currently, this alignment is typically performed using semi-automatic methods, which, while effective to some extent, are prone to errors that demand manual correction. Point cloud correspondence-based registration methods are promising to serve as a fully automatic solution. However, they may struggle in scenarios with limited intraoperative surface visibility, a common challenge in liver surgery, particularly in laparoscopic procedures, which we refer to as complete-to-partial ambiguity. We first illustrate this ambiguity by evaluating the performance of state-of-the-art learning-based point cloud registration methods on our carefully constructed in silico and in vitro datasets. Then, we propose a patches-to-partial matching strategy as a plug-and-play module to resolve the ambiguity, which can be seamlessly integrated into learning-based registration methods without disrupting their end-to-end structure. It has proven effective and efficient in improving registration performance for cases with limited intraoperative visibility. The constructed benchmark and the proposed module establish a solid foundation for advancing applications of point cloud correspondence-based registration methods in image-guided liver surgery. Code and datasets will be released.
@@ I Introduction
In liver surgery, preoperative imaging techniques such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) provide detailed information about critical anatomical structures, such as blood vessels and tumors. However, this information is not usually available during surgery due to the large footprint of the equipment, its immobility, and its prohibitively high cost. Consequently, intraoperative imaging is most commonly limited to lower resolution or partial information covering only a portion of the visible organ from modalities including intraoperative ultrasound [1, 2], cone-beam CT (CBCT) [3, 4], and optical cameras [5, 6, 7, 8]. To inform the surgeon of subsurface information from preoperative CT/MRI during surgery, a registration process is involved to compute the transformation between the preoperative and intraoperative data in image-guided surgery (IGS) systems, which is generally divided into two steps.
Step 1 is an initial rigid registration to resolve the problem that preoperative and intraoperative data are acquired in different coordinate systems. However, current IGS systems often perform this alignment semi-automatically [9], bringing considerable time consumption and susceptibility to errors [10].
Step 2 involves non-rigid registration methods to estimate tissue deformations, which depend on an accurate initial rigid registration.
@@
As a further challenge, the smooth surface of the liver lacks strong visual key points that could assist with matching partial point clouds and complicates current approaches to accurately align the visible intraoperative portion with the complete preoperative liver model (i.e., see Fig. 1 (d), where the target point cloud has a limited visible surface area). Besides the confusion associated with matching to the top or bottom of the liver, there is also ambiguity when registering a small point cloud to regions to regions adjacent to the liver dome.
As such, we argue that current learning-based point cloud registration methods are insufficient to handle the complete-to-partial ambiguity that frequently arises in LLS.
@@ I-A Contribution
In the P2P, we first select the most likely visible points from the source point cloud. Then, we generate several candidate patches of approximately the same size as the target point cloud. Finally, we match the target point cloud to these candidate patches and aggregate a robust, rigid point cloud transformation by selecting the best estimate across candidates. This module only involves feature resampling and rematching, allowing it to be parallelized for fast execution. Additionally, it is fully differentiable, making it suitable for integration into different correspondence-based point cloud and learning-based registration pipelines.
@@
We experimentally demonstrate that the current state-of-the-art correspondence-based registration methods are not sufficient when dealing with complete-to-partial ambiguity. We then show the effectiveness of our proposed module in dealing with low-visibility cases after seamless plug-and-play integration into two promising methods. Without any adjustment to loss functions, additional learnable modules, outlier rejection methods, etc., our approach significantly improves complete-to-partial registration.
@@ II-B Learning-based point cloud registration
Several learning-based methods [30, 31, 32] have been proposed for 3D-3D liver registration, yielding promising results. However, these outcomes should be interpreted with caution. For example, the simulation dataset used in [31] assumes one-to-one correspondence, as both the source and target point clouds are derived from the same mesh without post-processing. Consequently, models pre-trained or fine-tuned on these datasets may struggle when applied to real-world data. Furthermore, Guan et al. [30] and Zhang et al. [32] treat the intraoperative CT surface as the preoperative surface in the porcine dataset [33], resulting in source and target point clouds with no deformation.
In contrast, we carefully post-process our simulations to include soft tissue deformations, eliminate one-to-one correspondences, and remove rigid components from the simulated deformations. We evaluate the selected registration methods on both our simulation and phantom datasets, ensuring a more rigorous and realistic assessment.
@@ III Methodology
Figure 2: Illustration of the general paradigm of learning-based correspondence registration methods (top) and our plug-and-play patches-to-partial (P2P) module (bottom). Our proposed module generates candidate patches, samples their point-wise features from a learning-based correspondence registration method, and performs feature matching and rigid transformation estimation for each patch with the target point cloud. Finally, a candidate selection rule determines the optimal rigid transformation. We present the case where the patch number $k$ is set to 5.
This section introduces our proposed patches-to-partial strategy as a plug-and-play module that can be seamlessly integrated into current learning-based correspondence registration methods. We begin by presenting the general paradigm of learning-based correspondence registration methods (see §III-A), followed by a detailed module explanation (see §III-B). The overall framework is illustrated in Fig. 2.
@@ III-A Preliminaries
III-B Transferring complete-to-partial into patches-to-partial matching
@@
This challenge is similar to a Jigsaw puzzle, where a player tries different candidate spaces to find the correct fit. Inspired by this analogy, we propose converting the complete-to-partial point cloud matching problem into a patches-to-partial (P2P) matching process. By generating several candidate patches from the source point cloud, each roughly the size of the target, and matching them to the target, we can naturally solve the complete-to-partial ambiguity. The questions thus become clear:  How do we efficiently generate candidate patches from the source point cloud?  How do we select the best transformation?
For consistency in the presentation, we use the point-wise features $\mathbf{x}^{\mathbf{S}}$ and $\mathbf{x}^{\mathbf{T}}$, along with the estimated transformation ${\mathbf{R}_{c},\mathbf{t}_{c}}$ from existing correspondence-based registration methods, as our starting point to address these questions and validate this approach.
III-B1 Visible Source Point Estimation
@@
To narrow down the candidate points, we first design a method to select the top candidates from the source point cloud based on a visibility score, which we define according to a score matrix following [40, 13] as $S=\mathbf{x}^{\mathbf{S}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{N\times M}$, where $S$ represents the cosine similarity between source and target features. We then convert this score matrix into visibility scores for the source point cloud by summing over the target dimension: $S_{vis}=\sum_{j=1}^{M}S_{i,j}$. Finally, we select the top $M$ source points with the highest visibility scores to form the visible source point cloud.
III-B2 Patch Candidate Generation
@@
III-B3 Patch Sampling and Feature Computation
@@
III-B4 Patch-to-Target Matching
Each patch candidate generated in the previous step is then matched to the target point cloud to establish correspondences via rigid transformation estimation. Taking one patch as an example, we first compute the score matrix $S_{k}=\mathbf{x}^{\mathbf{S}_{k}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{N\times M}$, and apply the dual-softmax operation [40, 13, 38]:
@@ III-B4 Patch-to-Target Matching
III-B5 Optimal Candidate Selection
@@ IV-A1 In silico phantom dataset
For each liver model, we generated $10$ deformations using the pipeline proposed in [18] via a finite element model and used its default parameters. Some simulated deformations feature rigid transformations instead of deformation when the randomly generated zero boundary condition is small. To eliminate rigid transformation, we further align the undeformed and deformed models with the SVD using the volumetric vertices as fiducial markers. After the alignment, the average root mean square error of the fiducial markers is $3.37$ mm, with a maximum value of $11.38$ mm, which is similar to the deformation range in Sparse Non-rigid Registration Challenge dataset [45].
@@
For the testing dataset, we generated five target point clouds for each deformation. Consequently, the testing dataset contained $121$ source point clouds with $6,050$ target point clouds.
@@ V-A1 Performance against Baselines
We begin by evaluating the performance of the baseline methods under varying visibility ratios, with the quantitative results summarized in Table I.
@@
We then examine the results obtained by Lepard and LiverMatch with our proposed module, alongside those from RoITr quantitatively and qualitatively in Table I and Fig. 5, respectively. At the lowest visibility ratio range (i.e., visibility $\in[0.2,0.3]$), the registration errors for Lepard decrease from $9.47$ mm to $6.73$ mm, and for LiverMatch, from $12.85$ mm to $8.74$ mm following the integration of our proposed module. These improvements are statistically significant.
@@ V-A2 Sensitivity to Noise
V-A3 Visual Assessment
We further visualize the results in Fig. 4. In the first row, we show a case with a high visibility ratio as baseline, where all methods produce reasonable results. However, in the second and last rows, when the visibility ratio is low, all methods struggle to align the source to the target point cloud properly. After incorporating our proposed module, both LiverMatch and Lepard are able to achieve reasonable results.

### v2

@@ Abstract
In image-guided liver surgery, the initial rigid alignment between preoperative and intraoperative data, often represented as point clouds, is crucial for providing sub-surface information from preoperative CT/MRI images to the surgeon during the procedure. Currently, this alignment is typically performed using semi-automatic methods, which, while effective to some extent, are prone to errors that demand manual correction. Alternatively, correspondence-based point cloud registration methods further offer a promising fully automatic solution. However, they may struggle in scenarios with limited intraoperative surface visibility, a common challenge in liver surgery, particularly in laparoscopic procedures, which we refer to as complete-to-partial ambiguity. We first illustrate this ambiguity by evaluating the performance of state-of-the-art learning-based point cloud registration methods on our carefully constructed in silico and in vitro datasets. Then, we propose a patches-to-partial matching strategy as a plug-and-play module to resolve the ambiguity, which can be seamlessly integrated into learning-based registration methods without disrupting their end-to-end structure. This approach effectively improves registration performance, especially in low-visibility conditions, reducing registration errors to 6.7 mm (-29%) in silico and 12.5 mm (-40%) in vitro, compared to state-of-the-art performance achieved by Lepard of 9.5 mm and 20.7 mm, respectively. The constructed benchmark and the proposed module establish a solid foundation for advancing applications of point cloud correspondence-based registration methods in image-guided liver surgery. Our code and datasets will be released at https://github.com/zixinyang9109/P2P.
@@ I Introduction
In liver surgery, preoperative imaging techniques such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) provide detailed information about critical anatomical structures, such as blood vessels and tumors. However, this information is typically unavailable during surgery due to the large footprint of the equipment, its immobility, and its prohibitively high cost. Consequently, intraoperative imaging is most commonly limited to lower resolution or partial information covering only a portion of the visible organ from modalities including intraoperative ultrasound [1, 2], cone-beam CT (CBCT) [3, 4], and optical cameras [5, 6, 7, 8]. To inform the surgeon of subsurface information from preoperative CT/MRI during surgery, a registration process is involved to compute the transformation between the preoperative and intraoperative data in image-guided surgery (IGS) systems, which is generally divided into two steps.
Step 1 is an initial rigid registration to resolve the problem that preoperative and intraoperative data are acquired in different coordinate systems. However, current IGS systems often perform this alignment semi-automatically [9, 10], bringing considerable time consumption and susceptibility to errors [10].
Step 2 involves non-rigid registration methods to estimate organ deformations, which depend on an accurate initial rigid registration.
@@
I. Insufficient intraoperative-oriented design. Most learning-based point cloud registration methods are designed for partial-to-partial cases, where point cloud pairs are captured from different view angles and have roughly similar scales. These pairs typically come from RGB-D cameras [12] (Fig. 1 (a)), simulated depth maps from animated sequences [13] (Fig. 1 (b)), or synthetic objects [15, 16] (Fig. 1 (c)). Those pairs often feature prominent structures, such as edges and corners from natural scenes.
@@
As a further challenge, the smooth surface of the liver lacks strong visual key points that could assist with matching partial point clouds and complicates current approaches to accurately align the visible intraoperative portion with the complete preoperative liver model (i.e., see Fig. 1 (d), where the target point cloud has a limited visible surface area). Besides the confusion associated with matching to the top or bottom of the liver, there is also ambiguity when registering a small point cloud to regions adjacent to the liver dome.
As such, we argue that current learning-based point cloud registration methods are insufficient to address the complete-to-partial ambiguity that frequently arises in laparoscopic liver surgery. Traditional ICP-based methods [18, 19], often regarded as alternative solutions, also face limitations, including the need for manual intervention and high computational cost. Meanwhile, handcrafted features [20] struggle with the challenges posed by smooth surfaces that lack strong edges or curvature and lead to inherent ambiguity in point cloud matching.
@@ I-A Contribution
In the P2P module, we first identify the most likely visible points from the source point cloud by selecting the top points with the highest visibility scores, computed by summing feature similarities (see §III-C1). Patch nodes are then generated from these points and subsequently used to create candidate patches, each containing the same number of points as the target point cloud (see §III-C2 and III-C3). Finally, we match the target point cloud to these candidate patches and aggregate a robust, rigid point cloud transformation by selecting the best estimate across candidates (see §III-C4 and III-C5).This module only involves feature resampling and rematching, allowing it to be parallelized for fast execution. Additionally, it is fully differentiable, making it suitable for integration into different correspondence-based point cloud and learning-based registration pipelines.
@@
We experimentally demonstrate that the current state-of-the-art correspondence-based registration methods are not sufficient when dealing with complete-to-partial ambiguity. We then show the effectiveness of our proposed module in dealing with low-visibility cases after seamless plug-and-play integration into two promising methods, LiverMatch[22] and Lepard [13] (for additional details on the selection of these methods, refer to §VI-A). Without any adjustment to loss functions, additional learnable modules, outlier rejection methods, etc., our approach significantly improves complete-to-partial registration.
@@ II-B Learning-based point cloud registration
3D-3D initial rigid registration for liver surgery often relies on manual adjustment [9, 10] or landmark-based semi-automatic alignment [30, 31]. However, manual adjustment is prone to user-dependent errors. At the same time, anatomical landmarks can be difficult and ambiguous to annotate due to occlusions from surrounding anatomical structures and the smooth surface of the liver during surgery.
@@
Several learning-based methods [33, 22, 34] have been proposed for 3D-3D liver registration, yielding promising results. However, these outcomes should be interpreted with caution. For example, the simulation dataset used in [22] assumes one-to-one correspondence, as both the source and target point clouds are derived from the same mesh without post-processing. Consequently, models pre-trained or fine-tuned on these datasets may struggle when applied to real-world data. Furthermore, Guan et al. [33] and Zhang et al. [34] used the intraoperative CT surface as the preoperative surface in the porcine dataset [35], resulting in source and target point clouds with no deformation.
In contrast, we carefully post-process our simulations to include soft tissue deformations, eliminate one-to-one correspondences, and remove rigid components from the simulated deformations (see §IV-A1). We evaluate the selected registration methods on both our simulation and phantom datasets, ensuring a more rigorous and realistic assessment.
@@ III Methodology
Figure 2: Illustration of the general paradigm of learning-based correspondence registration methods (top) and our plug-and-play P2P module (bottom). The input source and target point clouds are voxelized to maintain similar densities, ensuring a uniform and continuous representation of the liver surface regions (see §IV-B1 and §VI-B). Our proposed module generates candidate patches, samples their point-wise features from a learning-based correspondence registration method, and performs feature matching and rigid transformation estimation for each patch with the target point cloud. Finally, a candidate selection rule determines the optimal rigid transformation. We present the case where the patch number $k$ is set to 5. The patch consists of the same number of points as the target point cloud (see §III-B and §III-C3).
This section introduces our proposed P2P strategy as a plug-and-play module that can be seamlessly integrated into current learning-based correspondence registration methods. We begin by presenting the general paradigm of learning-based correspondence registration methods (see §III-A), followed by a detailed module explanation (see §III-B). The overall framework is illustrated in Fig. 2.
@@ III-A Preliminaries
III-B Transferring Complete-to-Partial into P2P Matching
@@
This challenge is similar to a Jigsaw puzzle, where a player tries different candidate spaces to find the correct fit. Inspired by this analogy, we propose converting the complete-to-partial point cloud matching problem into a P2P matching process. By generating several candidate patches from the source point cloud, each roughly the size of the target, and matching them to the target, we can naturally solve the complete-to-partial ambiguity. The theoretical formulation can be written as:
${\color[rgb]{0,0,0}(R_{k},t_{k})=\mathop{\arg\min}_{R,\,t,\,S_{k}\in\mathcal{S}}E(S_{k},T,R,t)}$ (1)
where the complete source point cloud $\mathbf{S}\in\mathbb{R}^{N\times 3}$ is represented by several patches $\mathbf{S}_{k}\in\mathbb{R}^{M\times 3}$ with the same number of points as the target point cloud $\mathbf{T}^{M\times 3}$. Here, $E$ is the error function to select the optimal $R,t$ to align the source to the target point cloud.
Setting the point number of patches serves as a reasonable approximation and simplification, given the prerequisites we further discussed in (see §VI-B).
The theoretical foundation of our approach is two-fold. First, in comparison to the complete $\mathbf{S}$, a subset $\mathbf{S}_{k}$ that closely resembles the shape of $\mathbf{T}$ is likely to contain distinctive local structures, thereby reducing matching ambiguity. Second, the resulting set of transformations, $\mathcal{T}=\{(R_{1},t_{1}),(R_{2},t_{2}),...,(R_{k},t_{k})\}$, introduces multiple hypotheses, increasing the likelihood of identifying the optimal transformation.
The first argument is based on the premise that the ambiguity in complete-to-partial matching is minimized in the limit of the source point cloud approach the same extent and scale as the target point cloud. If the appropriate patch is selected, the registration problem effectively simplifies to a complete-to-complete scenario. The second argument then follows naturally: as the number of patches increases and selection is appropriately guided, the probability of identifying the correct transformation correspondingly improves.
The questions thus become clear:  How do we efficiently generate candidate patches from the source point cloud?  How do we select the best transformation? For consistency in the presentation, we use the point-wise features $\mathbf{x}^{\mathbf{S}}$ and $\mathbf{x}^{\mathbf{T}}$, along with the estimated transformation ${\mathbf{R}_{c},\mathbf{t}_{c}}$ from existing correspondence-based registration methods, as our starting point to address these questions and validate this approach.
III-C P2P Module
III-C1 Visible Source Point Estimation
@@
To narrow down the candidate points, we first design a method to select the top candidates from the source point cloud based on a visibility score, which we define according to a score matrix following [42, 13] as $S=\mathbf{x}^{\mathbf{S}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{N\times M}$, where $S$ represents the cosine similarity between source and target features. We then convert this score matrix into visibility scores for the source point cloud by summing over the target dimension: $S_{vis}=\sum_{j=1}^{M}S_{i,j}$. Finally, we select the top $M$ source points with the highest visibility scores to form the visible source point cloud for the next step to generate patch nodes.
III-C2 Patch Candidate Generation
@@
III-C3 Patch Sampling and Feature Computation
@@
III-C4 Patch-to-Target Matching
Each patch candidate generated in the previous step is then matched to the target point cloud to establish correspondences via rigid transformation estimation. Taking one patch as an example, we first compute the score matrix $S_{k}=\mathbf{x}^{\mathbf{S}_{k}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{{\color[rgb]{0,0,0}M}\times M}$, and apply the dual-softmax operation [42, 13, 40]:
@@ III-C4 Patch-to-Target Matching
III-C5 Optimal Candidate Selection
@@
Figure 4: Properties of the testing set from the in silico phantom dataset: (a) Distribution of RMS-TRE across all sample pairs after rigid alignment using volumetric vertices correspondences to remove rigid components. (b) Distribution of max-TRE per sample pair after rigid alignment. (c) Initial RMS-TRE across samples after applying random rigid transformations, used to evaluate rigid registration methods. (d) Visibility ratio of the target point cloud across samples after random cropping.
@@ IV-A1 In silico phantom dataset
For each liver model, we generated $10$ deformations using the pipeline proposed in [21] via a finite element model (neo-Hookean model) and used its default parameters. Some simulated deformations feature rigid transformations instead of deformation when the randomly generated zero boundary condition is small. To eliminate rigid transformation, we further align the undeformed and deformed models with the SVD using the volumetric vertices as fiducial markers. After the alignment, the average root mean square error of the fiducial markers is $3.37$ mm, with a maximum value of $11.38$ mm, which is similar to the deformation range in Sparse Non-rigid Registration Challenge dataset [47].
@@
For the testing dataset, we generated five target point clouds for each deformation. Consequently, the testing dataset contained $121$ source point clouds with $6,050$ target point clouds. The deformation and visibility of the testing dataset are shown in Fig. 4.
@@ V-A In Silico Phantom Validation
TABLE III: Registration errors on the in silico phantom dataset at different deformation ranges, within the visibility ratio range [0.2, 0.3). The average RMS-TRE ± standard deviation is reported in millimeters.
Deformation range (RMS-TRE) $(0,6)$ $[6,12]$
Procrustes 2.82 $\pm$ 1.51 7.30 $\pm$ 1.25
LiverMatch [22] 12.54 $\pm$ 13.75 17.24 $\pm$ 15.93
LiverMatch [22] + proposed P2P 8.54 $\pm$ 11.16 12.54 $\pm$ 5.13
Lepard [13] 8.92 $\pm$ 14.66 15.50 $\pm$ 13.82
Lepard [13] + proposed P2P 6.19 $\pm$ 5.50 12.85 $\pm$ 7.24
@@ V-A1 Performance against Baselines
We begin by evaluating the performance of the baselines under varying visibility ratios, with the quantitative results summarized in Table I.
@@
We then examine the results obtained by Lepard and LiverMatch with our proposed module, alongside those from RoITr quantitatively and qualitatively in Table I and Fig. 6, respectively. At the lowest visibility ratio range (i.e., visibility $\in[0.2,0.3]$), the registration errors for Lepard decrease from $9.47$ mm to $6.73$ mm, and for LiverMatch, from $12.85$ mm to $8.74$ mm, following the integration of our proposed module. These improvements over baseline results are confirmed to be statistically significant ($p<0.05$), based on the statistical analysis method described in §IV-B5.
@@ V-A2 Sensitivity to Noise
V-A3 Sensitivity to Deformation
Since rigid registration methods are expected to provide a reasonable initialization for subsequent non-rigid registration, we also examine their robustness under varying deformation levels and evaluate the effectiveness of the proposed P2P module. As shown in Table III, the P2P module consistently enhances registration performance across different deformation ranges within the low visibility ratio $[0.2,0.3)$. Specifically, for mild deformation $[0,6)$, registration errors are reduced from 12.54 mm to 8.54 mm for LiverMatch and from 8.92 mm to 6.19 mm for Lepard. Under more severe deformation $[6,12]$, errors decrease from 17.24 mm to 12.54 mm for LiverMatch and from 15.50 mm to 12.85 mm for Lepard. These results demonstrate the robustness of the proposed P2P strategy in handling both low visibility and large deformation scenarios.
V-A4 Success Rate
To determine the success rate of this automatic initial rigid registration approach, we define the success rate as the percentage of examples for which the RMS-TRE is below a certain threshold $\tau$.
In Fig. 7, we further demonstrate the effectiveness of the P2P module by comparing the success rate curves of LiverMatch and Lepard, both with and without the integration of P2P. Success rates achieved with P2P are always higher than results without P2P.
We also evaluate robustness using the success rate ($\tau=20$ mm) under different deformation levels. As shown in Table IV, the P2P module increases the success rate by 11.2% for LiverMatch and by 4.2% for Lepard under mild deformation (0-6 mm) and by 8.3% and 5.0%, respectively, under more severe deformation (6-12 mm).
These results further confirm the robustness and effectiveness of P2P in enhancing registration performance under challenging conditions.
Figure 7: Comparison of success rates at different thresholds ($\tau$) on the in silico phantom dataset within the visibility ratio range [0.2, 0.3).
TABLE IV: Registration success rate ($\tau=20\ mm$) in percentage on the in silico phantom dataset at different deformation ranges measured by RMS-TRE in millimeters, within the visibility ratio range [0.2, 0.3).
Deformation range (RMS-TRE) $(0,6)$ $[6,12]$
Procrustes 100.00 100.00
LiverMatch [22] 84.38 83.33
LiverMatch [22] + proposed P2P 95.54 91.67
Lepard [13] 94.49 86.67
Lepard [13] + proposed P2P 98.66 91.67
V-A5 Visual Assessment
We further visualize the results in Fig. 5. In the first row, we show cases with a high visibility ratio, where all methods produce reasonable results. However, in the second and last rows, when the visibility ratio is low, all methods struggle to align the source to the target point cloud properly. After incorporating our proposed module, both LiverMatch and Lepard achieved reasonable results.
Figure 8: Sensitivity study on patch number K, performed on the in vitro phantom dataset at different visibility ratios, using Lepard as the baseline for the proposed module.
TABLE V: Ablation study on the selection method and candidate proposal, performed on the in vitro phantom dataset at a visibility range of [0.2, 0.3), using Lepard as the baseline for the proposed module. The average RMS-TRE ± standard deviation is reported in millimeters.
Selection rule Inlier number 14.14 $\pm$ 16.70
Closest point distance (default) 12.45 $\pm$ 15.72
Candidate proposal w/o 14.91 $\pm$ 19.06
w/ (default) 12.45 $\pm$ 15.72
V-A6 Running Time Analysis
: Table VI summarizes the average running time of each registration method on the in silico phantom dataset. Traditional optimization-based GO-ICP is significantly slower than learning-based approaches, with an average time of 89.62 seconds per case. In contrast, learning-based methods complete registration within fractions of a second. The P2P module introduces only a slight increase in runtime while providing improved robustness to partial visibility, as previously demonstrated in Table I.
TABLE VI: Running time comparison of different registration methods on the In Silico phantom dataset.
Method Averaged Running Time (s)
GO-ICP 89.62
RoITr 0.25
RegTr 0.12
LiverMatch 0.07
LiverMatch + P2P 0.10
Lepard 0.10
Lepard + P2P 0.12